# Intelligent LiDAR Navigation: Leveraging External Information and Semantic Maps with LLM as Copilot*

### v2

@@ I INTRODUCTION
In real-world scenarios, robots operating in large dynamic environments face significant challenges, as their surroundings not always can be sensed or updated instantly. An illustrative, real-life example is a campus delivery robot encountering a path obstruction due to an unanticipated pipe repair closure, depicted in Fig. 1. Despite prior notice on public websites, the robot’s navigation algorithm remained unaware of this obstacle. With the capabilities of LLMs, which can process general text-based information, we now have the opportunity to integrate human-like contextual understanding into robotic systems. This way, in the above example, the robot could, by reading the website, from the beginning plan a path that avoids the construction areas, thus achieving a faster and shorter delivery route.
@@
For the scenario of this paper we argue, that in robot navigation, we can utilize an existing map that represents the environment and has semantic information. This map, in our case in the osmAG format [5], can be created by an Simultaneous Localization and Mapping (SLAM) approach or from CAD files. Utilizing LiDAR sensors we can localize the robot in this map and employ A* for optimal path planning once the goal coordinates are known. In our paper we utilize two LLM modules for two distinct tasks: 1) For human-robot-interaction we need to interpret the given language command to identify the navigation goal. For that the LLM leverages the semantic information encoded in the map to find goal coordinates, e.g. by finding the right room for “Please bring this item to the robotics training lab”. This module also utilized experience data to potentially adjust map costs. 2) We interpret external information to adjust map costs.
Figure 1: The figure above depicts a real-life situation encountered by a 3rd-party delivery robot on our University campus, where it is blocked by an intersection closure. Below the e-mail sent by Office of General Services announcing this closure is shown.
@@ III-C Navigation Event Tracking and Path Validation
As shown in Fig. 2, the NavigationEventMonitor is a second LLM that has two main tasks: It tracks events that may impact the robot’s future navigation tasks, such as elevator maintenance notifications. This is done once at the beginning of a navigation task. Additionally, this module module validates paths generated by the osmAGPathPlanner against tracked events (closed areas).
@@
The prompt used in this module to assess the path is shown in Fig. 4. It describes the task and specifies the input and output formats, without examples or chain-of-thought reasoning. The output of this module includes validation of the shortest path and, if the path is invalid, identifies which areas should be avoided and puts those in the ‘areas to avoid’ field, denoted as $A_{\text{invalid}}$. Since the LLM can infer potential impacts, for instance, if a notification states that main corridors will be closed for classroom renovations, the LLM may infer that other corridors in the same sector might also be affected. An additional field called ‘areas try to avoid’ denoted as $A_{soft}$, is included, assigning extra costs $d_{soft}$ to try to avoid these areas without avoiding them completely.
@@ IV-A1 Robot and World
Figure 5: This figure presents the paths from four experimental cases. For clarity, only the paths from our method and move_base with clearing between trials are displayed on the map. In the map, blue lines represent areas in the osmAG, while red lines indicate passages between them. Each case was tested in a Gazebo world with different door configurations (open or closed). Grey shaded zones represent restricted areas caused by events in case 1): a graduation party in the lobby, 2) classroom renovation in the D sector, 3) a fire drill in the lobby and B sector, and 4) a wireless access upgrade in the lobby and B sector. Blue shaded zones represent destination areas of each case. The inset figures show the path lengths for each trial, comparing our method, our method with ablation testing, move_base with and without obstacle layer clearing. The black box highlights where move_base failed, and the data within the yellow circle shows where the path intruded into restricted areas. In Case 4, only the path with NavigationEventMonitor avoids intruding into restricted areas.
@@ V-A Effectiveness of NavigationEventMonitor
The path validation results are summarized in Table I and evaluated by hand in terms of accuracy, recall, and false positive rate. Across 37 trials, 95 path approval requests were sent to ChatGPT-4o: 57 of 61 valid paths were approved (with is_valid set to ’true’ or the areas to avoid not in the path), and all 34 invalid paths were correctly flagged by setting is_valid to false. However, 14 requests out of 95, the responses were semantically correct but deviated from the specified output format, such as returning ’B sector’ or a room number instead of the exact area name, such that our navigation module could not parse them correctly. For DeepSeek-V3, 40 path approval requests were sent, evenly split between valid and invalid paths. All responses adhered strictly to the required output format. The findings highlight the high performance of both models in path assessment, confirming their practical utility in real-world applications.
TABLE I: Performance Comparison of ChatGPT-4o and DeepSeek-V3 in Path Validation (%)
Model Accuracy Recall False Positive Rate
ChatGPT-4o 0.96 1.0 0.93
DeepSeek-V3 0.93 0.9 0.95

### v3

@@ I INTRODUCTION
In real-world scenarios, robots operating in large dynamic environments face significant challenges, as their surroundings can not always be sensed or updated instantly. An illustrative, real-life example is a campus delivery robot encountering a path obstruction due to an unanticipated pipe repair closure, depicted in Fig. 1. Despite prior notice on public websites, the robot’s navigation algorithm remained unaware of this obstacle. With the capabilities of LLMs, which can process general text-based information, we now have the opportunity to integrate human-like contextual understanding into robotic systems. This way, in the above example, the robot could, by reading the website, from the beginning plan a path that avoids the construction areas, thus achieving a faster and shorter delivery route.
@@
For the scenario of this paper we argue, that in robot navigation, we can utilize an existing map that represents the environment and has semantic information. This map, in our case in the osmAG format [5], can be created by a Simultaneous Localization and Mapping (SLAM) approach or from CAD files. Utilizing LiDAR sensors we can localize the robot in this map and employ A* for optimal path planning once the goal coordinates are known. In our paper we utilize two LLM modules for two distinct tasks: 1) For human-robot-interaction we need to interpret the given language command to identify the navigation goal. For that the LLM leverages the semantic information encoded in the map to find goal coordinates, e.g. by finding the right room for “Please bring this item to the robotics training lab”. This module also utilized experience data to potentially adjust map costs. 2) We interpret external information to adjust map costs.
Figure 1: The figure above depicts a real-life situation encountered by a 3rd-party delivery robot on our University campus, where it is blocked by an intersection closure. Below the e-mail sent by Office of General Services announcing this closure is shown. Reprinted from [6].
@@ III-C Navigation Event Tracking and Path Validation
As shown in Fig. 2, the NavigationEventMonitor is a second LLM that has two main tasks: It tracks events that may impact the robot’s future navigation tasks, such as elevator maintenance notifications. This is done once at the beginning of a navigation task. Additionally, this module validates paths generated by the osmAGPathPlanner against tracked events (closed areas).
@@
The prompt used in this module to assess the path is shown in Fig. 4. It describes the task and specifies the input and output formats, without examples or chain-of-thought reasoning. The output of this module includes validation of the shortest path and, if the path is invalid, identifies which areas should be avoided and puts those in the ‘areas to avoid’ field, denoted as $A_{\text{invalid}}$. Since the LLM can infer potential impacts, for instance, if a notification states that main corridors will be closed for classroom renovations, the LLM may infer that other corridors in the same sector might also be affected. An additional field called ‘areas try to avoid’ denoted as $A_{soft}$, is included, assigning extra costs $d_{soft}$ to try to avoid these areas without avoiding them completely. Compared to pure A*, this module brings in additional computation time due to LLM inference, each online query taking on average about 2 seconds. However, by identifying and avoiding invalid or unfavorable areas, it ultimately reduces overall navigation time that would otherwise be spent traversing unsuitable paths.
@@
Utilizing osmAG’s room-level topology as the underlying graph structure, instead of the high-resolution occupancy grid, substantially decreases the computational burden of A* search and allows for more efficient plan and re-plan.
@@ IV-A1 Robot and World
Figure 5: This figure presents the paths from four experimental cases. For clarity, only the paths from our method (red) and move_base with clearing between trials (green) are displayed on the map. In the map, blue lines represent areas in the osmAG, while red lines indicate passages between them. Each case was tested in a Gazebo world with different door configurations (open or closed). Grey shaded zones represent restricted areas caused by events in case 1): a graduation party in the lobby, 2) classroom renovation in the D sector, 3) a fire drill in the lobby and B sector, and 4) a wireless access upgrade in the lobby and B sector. Blue shaded zones represent destination areas of each case. The inset figures show the path lengths for each trial, comparing our method, our method with ablation testing, move_base with and without obstacle layer clearing. The black box highlights where move_base failed, and the data within the yellow circle shows where the path intruded into restricted areas. In Case 4, only the path with NavigationEventMonitor avoids intruding into restricted areas.
@@ V-A Effectiveness of NavigationEventMonitor
The path validation results are summarized in Table I and evaluated by hand in terms of accuracy, recall, and false positive rate. Across 37 trials, 95 path approval requests were sent to ChatGPT-4o: 57 of 61 valid paths were approved (with is_valid set to ’true’ or the areas to avoid not in the path), and all 34 invalid paths were correctly flagged by setting is_valid to false. However, 14 requests out of 95, the responses were semantically correct but deviated from the specified output format, such as returning ‘B sector’ or a room number instead of the exact area name, such that our navigation module could not parse them correctly. For DeepSeek-V3, 40 path approval requests were sent, evenly split between valid and invalid paths. All responses adhered strictly to the required output format. The findings highlight the high performance of both models in path assessment, confirming their practical utility in real-world applications.
TABLE I: Path Validation Comparison: ChatGPT-4o vs. DeepSeek-V3 (%)
Model Accuracy Recall
ChatGPT-4o 0.96 1.0
DeepSeek-V3 0.93 0.9